# MoTM: Towards a Foundation Model for Time Series Imputation based on Continuous Modeling

### v2

@@ Abstract
Recent years have witnessed a growing interest for time series foundation models, with a strong emphasis on the forecasting task. Yet, the crucial task of out-of-domain imputation of missing values remains largely underexplored. We propose a first step to fill this gap by leveraging implicit neural representations (INRs). INRs model time series as continuous functions and naturally handle various missing data scenarios and sampling rates. While they have shown strong performance within specific distributions, they struggle under distribution shifts. To address this, we introduce MoTM (Mixture of Timeflow Models), a step toward a foundation model for time series imputation. Building on the idea that a new time series is a mixture of previously seen patterns, MoTM combines a basis of INRs, each trained independently on a distinct family of time series, with a ridge regressor that adapts to the observed context at inference. We demonstrate robust in-domain and out-of-domain generalization across diverse imputation scenarios (e.g., block and pointwise missingness, variable sampling rates), paving the way for adaptable foundation imputation models.
Keywords:
Imputation Foundation model Implicit Neural Representations.
@@ 1 Introduction
Oral presentation at the 10th Workshop on Advanced Analytics and Learning on Temporal Data (AALTD), ECML 2025
@@
  • We propose MoTM, a unified model capable of (i) handling various patterns of missing values at inference; (ii) achieving strong performance on out-of-domain datasets without retraining; (iii) effectively managing datasets sampled at different rates (e.g., 10min, 30min, 1h, 2h) by leveraging shared temporal structures across distributions. To the best of our knowledge, MoTM is the first model to meet all of these conditions (see Table 1).
@@ 2 Related Work
Recent advances in deep learning have led to an increasing number of models for time series imputation. BRITS [4] pioneered the use of bidirectional RNNs for imputation, while subsequent methods explored alternative architectures such as GANs [18, 19, 17], VAEs [10], diffusion models [25], matrix factorization techniques [16], and Transformer-based models like SAITS [9]. Despite their success, these models assume regularly sampled data limiting their flexibility in real-world applications. Moreover, their generalization capabilities in out-of-domain (OOD) settings remain limited.
@@ 3.1 Problem Setting
A visual illustration of the notations is provided below in Fig. 1.
@@ 3.2 Key Components
In contrast to plain INRs, which are typically designed to represent a single function (e.g., one time series), TimeFlow is a generalizable INR, able to model an entire collection $(x^{(j)})_{j}$ of time series. This is achieved by incorporating per-sample modulations (through per-sample additive bias $\psi^{(j)}$) that condition the function $f_{\theta}$ on each specific instance. See Fig. 2 for a visualization. We refer to the original TimeFlow paper for a detailed description of this mechanism [14].
@@
where $\mathbf{W}^{(j)}\in\mathbb{R}^{(N_{\text{train}}\cdot d+1)\times 1}$ contains the regression coefficients and the intercept. A visualization of this process is shown in Fig. 3.
@@ 4.1 Experiments on Synthetic Data
We generate three synthetic datasets summarized in Table 2, using Gaussian Processes (GP) via KernelSynth [1]. Each synthetic series is the sum of three components, namely: (i) a smooth trend sampled from a GP with RBF kernel; (ii) a seasonal component sampled from a GP with a periodic Exponential-Sine-Squared kernel; (iii) a residual sampled from a GP with white noise kernel. We create $N_{\texttt{train}}=2$ datasets for pretraining, ks1D (resp., ks1W) with an hourly sampling frequency (resp., half-hourly) and daily (resp., weekly) periodicities. A chronological 0.75 - 0.25 train - test split is applied on both ks1D and ks1W. The third dataset ks1D1W, sampled at a 15-min rate with both a daily and a weekly periodic component, is kept for inference only.
@@
Following the procedure described in Section 3, we train one TimeFlow model on ks1D and another on ks1W. Both models are trained on time grids spanning four-week periods. To condition each TimeFlow on partially observed time series, we simulate missing data during training by randomly removing a subset of timesteps from the grids $\mathcal{T}_{obs}$ at each optimization step, resulting in sparse inputs with a missingness ratio $\tau\in\{0.01,0.2,0.3,\dots,0.7\}$. At inference time, as detailed in Section 3, MoTM reuses both pretrained TimeFlow models to fit a ridge regressor on new partially observed time series.
@@
We compare MoTM against several baselines: $\bullet$ TimeFlow 1D and TimeFlow 1W respectively use predictions from a single TimeFlow model trained on ks1D and ks1W $\bullet$ Linear performs standard linear interpolation between observed points $\bullet$ Repeat imputes each missing value by copying the most recent available observation from the desired seasonnality. Specifically, it uses the last available value from the previous day in the case of daily seasonality, or from the previous week in the case of weekly seasonality. $\bullet$ Mixture I aggregates the predictions of both pretrained TimeFlow models using a softmax-weighted average, with weights derived from the negative reconstruction scores on the available context $\mathcal{T}^{j}_{obs}$bullet Mixture II fits a ridge regressor on the observed context $\mathcal{T}^{j}_{obs}$, using the output predictions of both pretrained TimeFlow models as features.
@@ Results.
Table 3 shows that MoTM achieves good performances on the in-domain datasets ks1D and ks1W, on a par with the strong supervised baselines. We also note that Mixture II, the ridge regression on top of the TimeFlows predictions, emerges as a simple method to match the best pretrained model at inference.
@@
Visually, Fig. 4 showcases MoTM’s ability to fit the context and impute missing values in challenging scenarios.
@@ 4.2 Experiments on Real World Data
For these experiments, the MoTM model is built on $N_{\text{train}}=3$ TimeFlow models, each trained on one of the following datasets: (i) Electricity, (ii) Solar, (iii) and SpanishW-T. These datasets were selected to form a rich training domain, with diversity in sampling frequencies, seasonal patterns (daily and weekly for Electricity, daily for Solar and SpanishW-T), sample sizes (ranging from 5 to 321), and within-series variability. For out-of-domain evaluation, we use standard benchmarks: Traffic, ETTh1, ETTh2, Weather, and Spanish Energy. Their key characteristics are summarized in Table 4.
@@
[…]
@@
The hyperparameters for TimeFlow are the same as in Section 4.1, except for an INR hidden size of 256. We allow training for 40k epochs on Electricity and Solar, 20k epochs on SpanishW-T. For MoTM, the grid search yields $\lambda=0.5$ for pointwise imputation and $\lambda=1$ for block imputation.
@@ Results.
Several observations emerge from the test results reported in Table 5. (i) MoTM consistently outperforms the zero-shot baseline MOMENT, which performs poorly across all datasets, highlighting the limitations of foundation models without adaptation. (ii) MoTM achieves substantial gains over the supervised TimeFlow model on both in-domain (ID) and out-of-domain (OOD) datasets, demonstrating the benefits of multi-source training and the effectiveness of ridge-based adaptation. (iii) Compared to other supervised deep learning methods, MoTM delivers competitive or superior performance — particularly against BRITS, reducing its error by $30.6\%$ on ID and $24.0\%$ on OOD datasets. In addition, MoTM slightly outperforms SAITS in the ID setting ($12.6\%$ improvement on average), while remaining competitive in the OOD setting, even surpassing SAITS on datasets like ETTh1 or Weather in the Block scenarios. (iv) MoTM also clearly outperforms statistical baselines such as linear interpolation and value repetition, with an average relative improvement of $22.5\%$ and $32.2\%$ in-domain, and $13.2\%$ and $31.1\%$ out-of-distribution.
Overall, the results highlight the benefits of continuous modeling for time series imputation, as well as the effectiveness of MoTM’s simple adaptation strategy to generalize to new domains. Qualitatively, Fig. 5 illustrates a block imputation example from the OOD dataset SpanishE. Visually, MoTM demonstrates strong imputation capabilities on this complex sample.
@@
Can the performance of MoTM be matched by a single TimeFlow with ridge adaptation at inference? To elucidate this question, we perform an ablation on $N_{\texttt{train}}$. Fig. 6 shows the evolution of the test MAE on four datasets as the mixture grows from one component with ridge adaptation (Electricity) to all three components. In general, increasing $N_{\texttt{train}}$ does improve the test metrics on ID and OOD datasets, showing that MoTM leverages its multi-source pretraining. However, certain datasets and settings such as block imputation on Traffic, do not benefit from more base components. This calls for further improvement of the orchestration mechanism, e.g. through a better tuning of $\lambda$.
@@
At first glance, the inference procedure of MoTM may appear computationally expensive, due to the combination of adapting TimeFlow models and fitting the ridge regression. To assess its practical cost, we report in Table 6 the inference time of MoTM on the largest test dataset, Traffic. On a single H100 GPU, imputing 83k segments of length 672 takes approximately 61 seconds in total, corresponding to roughly $0.7$ milliseconds per segment. This suggests that MoTM remains computationally efficient at inference time, even on large-scale data. For comparison, SAITS requires full retraining on this out-of-distribution dataset and takes approximately 3h16 to reach the performance reported in Table 5 on the same task.
@@ 5 Conclusion and Discussion
While MoTM inference is slightly slower than single-model approaches due to its mixture structure, it remains efficient relative to the substantial training time required by comparable models retrained from scratch, as shown in Table 6. Overall, the performance gains brought by MoTM are not uniform across all ID datasets. For instance, the Solar dataset shows limited improvement, highlighting that the benefit of model mixing depends on the underlying data distribution. Identifying the most effective combination of models remains thus an open challenge. Moreover, the supervised baseline SAITS outperforms MoTM in several settings, suggesting that there is still room for improvement. Future work will explore the construction of large, unified databases enabling the training of TimeFlow models at scale for diverse distributions, and the integration of more expressive orchestration mechanisms to further enhance inference quality.

### v3

@@ Abstract
Recent years have witnessed a growing interest for time series foundation models, with a strong emphasis on the forecasting task. Yet, the crucial task of out-of-domain imputation of missing values remains largely underexplored. We propose a first step to fill this gap by leveraging implicit neural representations (INRs). INRs model time series as continuous functions and naturally handle various missing data scenarios and sampling rates. While they have shown strong performance within specific distributions, they struggle under distribution shifts. To address this, we introduce MoTM (Mixture of Timeflow Models), a step toward a foundation model for time series imputation. Building on the idea that a new time series is a mixture of previously seen patterns, MoTM combines a basis of INRs, each trained independently on a distinct family of time series, with a ridge regressor that adapts to the observed context at inference. We demonstrate robust in-domain and out-of-domain generalization across diverse imputation scenarios (e.g., block and pointwise missingness, variable sampling rates), paving the way for adaptable foundation imputation models. https://github.com/EtienneLnr/MoTM
@@ 1 Introduction
10th Workshop on Advanced Analytics and Learning on Temporal Data (AALTD), ECML 2025
@@
  • We propose MoTM, a unified model capable of (i) handling various patterns of missing values at inference; (ii) achieving strong performance on out-of-domain datasets without retraining; (iii) effectively managing datasets sampled at different rates (e.g., 10min, 30min, 1h, 2h) by leveraging shared temporal structures across distributions. To the best of our knowledge, MoTM is the first model to meet all of these conditions (see Table˜1).
@@ 2 Related Work
Recent advances in deep learning have led to an increasing number of models for time series imputation. BRITS [4] pioneered the use of bidirectional RNNs for imputation, while subsequent methods explored alternative architectures such as GANs [17, 18], VAEs [10], diffusion models [24], matrix factorization techniques [16], and Transformer-based models like SAITS [9]. Despite their success, these models assume regularly sampled data limiting their flexibility in real-world applications. Moreover, their generalization capabilities in out-of-domain (OOD) settings remain limited.
@@ 3.1 Problem Setting
A visual illustration of the notations is provided below in Fig.˜1.
@@ 3.2 Key Components
In contrast to plain INRs, which are typically designed to represent a single function (e.g., one time series), TimeFlow is a generalizable INR, able to model an entire collection $(x^{(j)})_{j}$ of time series. This is achieved by incorporating per-sample modulations (through per-sample additive bias $\psi^{(j)}$) that condition the function $f_{\theta}$ on each specific instance. See Fig.˜2 for a visualization. We refer to the original TimeFlow paper for a detailed description of this mechanism [14].
@@
where $\mathbf{W}^{(j)}\in\mathbb{R}^{(N_{\text{train}}\cdot d+1)\times 1}$ contains the regression coefficients and the intercept. A visualization of this process is shown in Fig.˜3.
@@ 4.1 Experiments on Synthetic Data
We generate three synthetic datasets summarized in Table˜2, using Gaussian Processes (GP) via KernelSynth [1]. Each synthetic series is the sum of three components, namely: (i) a smooth trend sampled from a GP with RBF kernel; (ii) a seasonal component sampled from a GP with a periodic Exponential-Sine-Squared kernel; (iii) a residual sampled from a GP with white noise kernel. We create $N_{\texttt{train}}=2$ datasets for pretraining, ks1D (resp., ks1W) with an hourly sampling frequency (resp., half-hourly) and daily (resp., weekly) periodicities. A chronological 0.75 - 0.25 train - test split is applied on both ks1D and ks1W. The third dataset ks1D1W, sampled at a 15-min rate with both a daily and a weekly periodic component, is kept for inference only.
@@
Following the procedure described in Section˜3, we train one TimeFlow model on ks1D and another on ks1W. Both models are trained on time grids spanning four-week periods. To condition each TimeFlow on partially observed time series, we simulate missing data during training by randomly removing a subset of timesteps from the grids $\mathcal{T}_{obs}$ at each optimization step, resulting in sparse inputs with a missingness ratio $\tau\in\{0.01,0.2,0.3,\dots,0.7\}$. At inference time, as detailed in Section˜3, MoTM reuses both pretrained TimeFlow models to fit a ridge regressor on new partially observed time series.
@@
We compare MoTM against several baselines: $\bullet$ TimeFlow 1D and TimeFlow 1W respectively use predictions from a single TimeFlow model trained on ks1D and ks1W $\bullet$ Linear performs standard linear interpolation between observed points $\bullet$ Repeat imputes each missing value by copying the most recent available observation from the desired seasonnality. Specifically, it uses the last available value from the previous day in the case of daily seasonality, or from the previous week in the case of weekly seasonality. $\bullet$ Mixture I aggregates the predictions of both pretrained TimeFlow models using a softmax-weighted average, with weights derived from the negative reconstruction scores on the available context $\mathcal{T}^{j}_{obs}$ $\bullet$ Mixture II fits a ridge regressor on the observed context $\mathcal{T}^{j}_{obs}$, using the output predictions of both pretrained TimeFlow models as features.
@@ Results.
Table˜3 shows that MoTM achieves good performances on the in-domain datasets ks1D and ks1W, on a par with the strong supervised baselines. We also note that Mixture II, the ridge regression on top of the TimeFlows predictions, emerges as a simple method to match the best pretrained model at inference.
@@
Visually, Fig.˜4 showcases MoTM’s ability to fit the context and impute missing values in challenging scenarios.
@@ 4.2 Experiments on Real World Data
For these experiments, the MoTM model is built on $N_{\text{train}}=3$ TimeFlow models, each trained on one of the following datasets: (i) Electricity, (ii) Solar, (iii) and SpanishW-T. These datasets were selected to form a rich training domain, with diversity in sampling frequencies, seasonal patterns (daily and weekly for Electricity, daily for Solar and SpanishW-T), sample sizes (ranging from 5 to 321), and within-series variability. For out-of-domain evaluation, we use standard benchmarks: Traffic, ETTh1, ETTh2, Weather, and Spanish Energy. Their key characteristics are summarized in Table˜4.
@@
All datasets are split chronologically intro train, validation and test fractions with respective ratios 0.7 - 0.1 - 0.2. The ridge regularization coefficient $\lambda$ is selected from a grid search over $\{0.01,\,0.1,\,0.5,\,1,\,5,\,10\}$ on the validation split of the three training datasets. The rest of the protocol for training and evaluation follows the one described for synthetic data.
@@
The hyperparameters for TimeFlow are the same as in Section˜4.1, except for an INR hidden size of 256. We allow training for 40k epochs on Electricity and Solar, 20k epochs on SpanishW-T. For MoTM, the grid search yields $\lambda=0.5$ for pointwise imputation and $\lambda=1$ for block imputation.
@@ Results.
Several observations emerge from the test results reported in Table˜5. (i) MoTM consistently outperforms the zero-shot baseline MOMENT, which performs poorly across all datasets, highlighting the limitations of foundation models without adaptation. (ii) MoTM achieves substantial gains over the supervised TimeFlow model on both in-domain (ID) and out-of-domain (OOD) datasets, demonstrating the benefits of multi-source training and the effectiveness of ridge-based adaptation. (iii) Compared to other supervised deep learning methods, MoTM delivers competitive or superior performance — particularly against BRITS, reducing its error by $30.6\%$ on ID and $24.0\%$ on OOD datasets. In addition, MoTM slightly outperforms SAITS in the ID setting ($12.6\%$ improvement on average), while remaining competitive in the OOD setting, even surpassing SAITS on datasets like ETTh1 or Weather in the Block scenarios. (iv) MoTM also clearly outperforms statistical baselines such as linear interpolation and value repetition, with an average relative improvement of $22.5\%$ and $32.2\%$ in-domain, and $13.2\%$ and $31.1\%$ out-of-distribution.
Overall, the results highlight the benefits of continuous modeling for time series imputation, as well as the effectiveness of MoTM’s simple adaptation strategy to generalize to new domains. Qualitatively, Fig.˜5 illustrates a block imputation example from the OOD dataset SpanishE. Visually, MoTM demonstrates strong imputation capabilities on this complex sample.
@@
Can the performance of MoTM be matched by a single TimeFlow with ridge adaptation at inference? To elucidate this question, we perform an ablation on $N_{\texttt{train}}$. Fig.˜6 shows the evolution of the test MAE on four datasets as the mixture grows from one component with ridge adaptation (Electricity) to all three components. In general, increasing $N_{\texttt{train}}$ does improve the test metrics on ID and OOD datasets, showing that MoTM leverages its multi-source pretraining. However, certain datasets and settings such as block imputation on Traffic, do not benefit from more base components. This calls for further improvement of the orchestration mechanism, e.g. through a better tuning of $\lambda$.
@@
At first glance, the inference procedure of MoTM may appear computationally expensive, due to the combination of adapting TimeFlow models and fitting the ridge regression. To assess its practical cost, we report in Table˜6 the inference time of MoTM on the largest test dataset, Traffic. On a single H100 GPU, imputing 83k segments of length 672 takes approximately 61 seconds in total, corresponding to roughly $0.7$ milliseconds per segment. This suggests that MoTM remains computationally efficient at inference time, even on large-scale data. For comparison, SAITS requires full retraining on this out-of-distribution dataset and takes approximately 3h16 to reach the performance reported in Table˜5 on the same task.
@@ 5 Conclusion and Discussion
While MoTM inference is slightly slower than single-model approaches due to its mixture structure, it remains efficient relative to the substantial training time required by comparable models retrained from scratch, as shown in Table˜6. Overall, the performance gains brought by MoTM are not uniform across all ID datasets. For instance, the Solar dataset shows limited improvement, highlighting that the benefit of model mixing depends on the underlying data distribution. Identifying the most effective combination of models remains thus an open challenge. Moreover, the supervised baseline SAITS outperforms MoTM in several settings, suggesting that there is still room for improvement. Future work will explore the construction of large, unified databases enabling the training of TimeFlow models at scale for diverse distributions, and the integration of more expressive orchestration mechanisms to further enhance inference quality.
6 Reproducibility
All code and data used in this work are publicly available to ensure reproducibility of the results.